The age of father is twice the square of the age of his son . Eight years hence , father's age will be thrice times as that of his son . Find their present age ?

Answer by richwmiller(17219)

f=2s^2
f+8=3*(s+8)
f+8=3s+24
f=3s+16
2s^2=3s+16
2s^2-3s-16=0
positive answers only